calculate the mass of oxygen needed to react 10.0g of calcium to form calcium oxide


1
Expert's answer
2021-09-12T23:16:36-0400

Solution:

Calculate the moles of calcium (Ca):

The molar mass of Ca is 40 g/mol

Hence,

(10.0 g Ca) × (1 mol Ca / 40 g Ca) = 0.25 mol Ca


Balanced chemical equation:

2Ca + O2 → 2CaO

According to stoichiometry:

2 mol of Ca reacts with 1 mol of O2

Thus, 0.25 mol of Ca reacts with:

(0.25 mol Ca) × (1 mol O2 / 2 mol Ca) = 0.125 mol O2


Calculate the mass of oxygen (O2):

The molar mass of O2 is 32 g/mol

Hence,

(0.125 mol O2) × (32 g O2 / 1 mol O2) = 4 g O2


Answer: 4 grams of oxygen (O2).